OpenStudy (hannahwaller5):

Will Give Medal! Solve: 16x^2 − 81 = 0. x = 4, 9 x = −4, −9 x = 9 over 4, x = −9 over 4 x = 81 over 16, x = −81 over 16

OpenStudy (shamim):

Do uknow a^2-b^2=?

OpenStudy (shamim):

a^2-b^2=(a+b)(a-b)

OpenStudy (hannahwaller5):

what do you mean...?

OpenStudy (shamim):

U hv to usethis formula in math

OpenStudy (hannahwaller5):

oh ok i THINK i know what you are talking about...

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle 16x^2-81 =0 }\) \(\color{#000000}{ \displaystyle 16x^2=81 }\) \(\color{#000000}{ \displaystyle (4x)^2=9^2 }\) \(\color{#000000}{ \displaystyle 4x=\pm 9 }\)

OpenStudy (shamim):

Now tell me (4x)^2=?

OpenStudy (hannahwaller5):

Ok i got the 3rd one is that correct?

OpenStudy (hannahwaller5):

and16x

OpenStudy (shamim):

Yes correct!!!!

OpenStudy (hannahwaller5):

yay (:

OpenStudy (hannahwaller5):

Thanka

OpenStudy (shamim):

U r most welcome!!!!